OpenStudy (anonymous):

Divide -2x^3 - 4x^2 +3x +2 by x-3

OpenStudy (anonymous):

@littlebit can you help?

OpenStudy (anonymous):

sure i can try

OpenStudy (anonymous):

4 - 4x - 3x² = 0 x² + 4/3x = 4/3 x² + 2/3x = 4/3 + (2/3)² x² + 2/3x = 12/9 + 4/9 (x + 2/3)² = 16/9 x + 2/3 = 4/3 x = 4/3 - 2/3, x = 2/3 x = - 4/3 - 2/3, x = - 6/3, x = - 2 Answer: x = 2/3, - 2 Proof (x = 2/3): 4 - 4(2/3) = 3(2/3)² 4 - 8/3 = 3(4/9) 12/3 - 8/3 = 4/3 4/3 = 4/3 Proof (x = - 2): 4 - 4(- 2) = 3(- 2)² 4 + 8 = 3(4) 12 = 12

OpenStudy (anonymous):

it always have to do that grr anyways im not sure anyone want to check me?

OpenStudy (anonymous):

the answer choices i have are A. -2x^2-10x-27, R (-79) B. -2x^2-10x-27 C. -2x^2-10x-27, R 83 D. -2x^2-10x+27, R 83

OpenStudy (anonymous):

@littlebit do you know which of those is the answer?

OpenStudy (anonymous):

no im not sure sorry but im going to see if i can get anyone to help

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@lmw buddy can u help since your in 12th?

OpenStudy (anonymous):

@Dbzfan836 @sammixboo

OpenStudy (kamibug):

Use synthetic division since the divisor is a binomial with a degree of 1.  -2x^2-10x-27, R (-79)